A Ferris wheel of radius 10 meters starts from res
Answer #1

a) s = 0.5at^2

2 pi = 0.5*a*20^2

a = [2 pi / (200)] = 0.0314 rad/s^2 answer

b) w = 0.0314*20 = 0.628 rad/s answer

c) v = 0.628*10 = 6.28 m/s

d) a = v^2/r = 6.28^2/10 = 3.944 m/s^2